Consider the following balanced reaction: 3C(s) +
Answer #1

H2(g) + 1/2 O2 ............> H2O DH0 = - 285.8 KJ ....................1

C(S) + O2 .................> CO2    DH0 = - 393.5 KJ.......................2

C3H8 + 5O2 .............> 3CO2 + 4H2O DH0 = -2219.9 KJ..........................3

Do 3*eq.2 + 4*eq.1 - eq.3

4H2(g) + 2O2 ............> 4H2O DH0 = -1143.2 Kj

3C(S) +3O2 ...............> 3CO2   DH0 = - 1180.5 Kj

3CO2 + 4H2O .............> C3H8 + 5O2 DH0 = 2219.9 KJ

.....................................................................................................

3C(S) + 4H2 ..............>C3H8(g)   DH = (-1143.2 - 1180 + 2219.9) = -103.3 KJ